# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ C.3 Additional Lemmas
$\|\psi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\left(1+\frac{1}{2\sqrt{T}}C\mu\kappa^{2}r^{1/2}\sqrt{\frac{mn}{\left(m+n\right)\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$
Combining this result with Proposition D.2, we get that with probability larger than $1-\delta,$
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+C_{1}\sqrt{\log(e/\delta)}\mu^{2}\kappa^{2}r\omega_{\min}^{-1}\left(m\wedge n\right)^{-1/2}\frac{\log^{3/2}\left(\left(m+n\right)T/\delta\right)}{T}$
$+C_{2}\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
for some universal constants $C_{1},C_{2}>0$. Since $\sqrt{\log(e/\delta)}\mu^{2}\kappa^{2}r\omega_{\min}^{-1}\left(m\wedge n\right)^{-1/2}\lesssim\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\log^{3}\left(\left(m+n\right)T/\delta\right)$, adapting the previous arguments ensures that under the assumptions of Proposition D.2,
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+J\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$ (39)
with probability larger than $1-\delta.$
Transforming this result into a sample complexity guarantee as in the proof of Theorem 6.2 yields that $\hat{v}_{\text{\tt RS-PE}}$ is $\left(\varepsilon,\delta\right)$-PAC as soon as
$T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
for $K_{2}:=\max\left(\frac{1}{w_{\min}},\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\right).$
Recall that when $M$ and $\omega$ are homogeneous, $K_{2}=\Theta\left(mn\right)$. In this case, the above sample complexity guarantee can be reformulated as
$T$ $\gtrsim\frac{\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)$
@@
$\varepsilon\lesssim\frac{\left(m+n\right)}{mn}\frac{\log\left(e/\delta\right)}{\log^{3}\left(\left(m+n\right)T/\delta\right)}.$ (40)
@@ D.3 Additional Lemmas
A direct computation shows that the diagonal entries of $\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}$ can all be written as $\sum_{i,j}U^{2}_{i,k}V^{2}_{j,l}$, $\sum_{i,j}U^{2}_{\perp,i,k}V_{j,l}^{2}$ or $\sum_{i,j}U^{2}_{i,k}V_{\perp,j,l}^{2}$ for some $k$ and $l$. All of these terms are equal to $1$ by the orthogonality of $\begin{bmatrix}U&U_{\perp}\\ \end{bmatrix}$ and $\begin{bmatrix}V&V_{\perp}\\ \end{bmatrix}$. The entries outside of the diagonal can all be factorized by a scalar product between two columns of $\begin{bmatrix}U&U_{\perp}\\ \end{bmatrix}$ or two columns of $\begin{bmatrix}V&V_{\perp}\\ \end{bmatrix}$, which is equal to 0 by the same orthogonality argument. The same reasoning holds for $\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}$, as $\begin{bmatrix}\widehat{U}&\widehat{U}_{\perp}\\ \end{bmatrix}$ and $\begin{bmatrix}\widehat{V}&\widehat{V}_{\perp}\\ \end{bmatrix}$ are also orthogonal. The other part of the statement directly follows from the identities $\omega_{\max}\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}\succcurlyeq\Lambda\succcurlyeq\omega_{\min}\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}$ and $\omega_{\max}\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}\succcurlyeq\Lambda_{\phi}\succcurlyeq\omega_{\min}\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}$. ∎
@@
  • $\left|\|\phi_{\nu}\|^{2}_{2}-\|\psi_{\nu}\|^{2}_{2}\right|\leq 3\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)$
@@
$\phi_{i,j}^{\top}\phi_{i^{\prime},j^{\prime}}=\left(\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}\right)\left(\widehat{U}_{j}^{\top}\widehat{V}_{j^{\prime}}\right)+\left(\widehat{U}_{\perp,i}^{\top}\widehat{U}_{\perp,i^{\prime}}\right)\left(\widehat{U}_{j}^{\top}\widehat{V}_{j^{\prime}}\right)+\left(\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}\right)\left(\widehat{V}_{\perp,j}^{\top}\widehat{V}_{\perp,j^{\prime}}\right)$
@@
Note that by orthogonality, $\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}+\widehat{U}_{\perp,i}^{\top}\widehat{U}_{\perp,i^{\prime}}=U_{i}^{\top}U_{i^{\prime}}+U_{\perp,i}^{\top}U_{\perp,i^{\prime}}=0$ when $i\neq i^{\prime}$, and $\widehat{U}_{j}^{\top}\widehat{V}_{j^{\prime}}+\widehat{V}_{\perp,j}^{\top}\widehat{V}_{\perp,j^{\prime}}=V_{j}^{\top}V_{j^{\prime}}+V_{\perp,j}^{\top}V_{\perp,j^{\prime}}=0$ when $j\neq j^{\prime}$. Thus, two out of the three terms in the equalities above cancel out when $(i,j)\neq(i^{\prime},j^{\prime})$. When $(i,j)=(i^{\prime},j^{\prime})$, we can combine two terms by using the fact that $\|U_{i}\|^{2}_{2}+\|U_{\perp,i}\|^{2}_{2}=1$ to write $\psi_{i,j}^{\top}\psi_{i,j}=\|V_{j}\|^{2}_{2}+\|U_{i}\|^{2}_{2}\|V_{\perp,j}\|^{2}_{2}.$ In any case, each term in the difference can be controlled by the subspace recovery error. For example,
@@
$\leq 2\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right).$
@@
$\phi_{i,j}^{\top}\phi_{i^{\prime},j^{\prime}}-\psi_{i,j}^{\top}\psi_{i^{\prime},j^{\prime}}\leq\left(2+\mathds{1}_{(i,j)=(i^{\prime},j^{\prime})}\right)\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right).$
@@
$\leq\left(\sum_{(i,j),(i^{\prime},j^{\prime})}\nu_{i,j}\nu_{i^{\prime},j^{\prime}}\left(2+\mathds{1}_{(i,j)=(i^{\prime},j^{\prime})}\right)\right)\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)$
$=\left(2+\sum_{i,j}\nu_{i,j}^{2}\right)\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right),$
@@
There are two interesting things to mention about the second inequality of Lemma D.12. First, it is tight in the sense that it is attained for $r=m\wedge n$. Indeed, assume that $m\wedge n=n$ without loss of generality. On the one hand, $\|\psi_{i,j}\|^{2}_{2}=\|V_{j}\|^{2}_{2}+\|U_{i}\|^{2}_{2}\left(1-\|V_{j}\|^{2}_{2}\right)=1$ since $V_{j}$ is a row of the orthogonal matrix $V$. On the other hand, by orthogonality of $U$, $\mu=\|U\|_{2\to\infty}=1$ so that $\frac{\mu^{2}r\left(m+n-\mu^{2}r\right)}{mn}=1$. Second, it directly entails an upper bound on the norm of $\psi_{\pi}.$ Indeed,
$\|\psi_{\pi}\|_{2}$ $=\|\sum_{i,j}w^{\pi}_{i,j}\psi_{i,j}\|_{2}$
$\leq\sum_{i,j}w^{\pi}_{i,j}\|\psi_{i,j}\|_{2}$
$\leq\max_{i,j}\|\psi_{i,j}\|_{2}$
where we have used $\sum_{i,j}w^{\pi}_{i,j}=1$ in the last inequality. This ensures that
@@
In this section, we recall that we assume that $\xi_{t}\sim\mathcal{N}(0,1)$ and we choose the rank factorization $M=U\Sigma V^{\top}=PQ^{\top}$ with $P=U\sqrt{\Sigma}$ and $Q=V\sqrt{\Sigma}$. Furthermore, we denote by $N_{i}$ the $i$-th row vector of a matrix $N$. As noted in (Jun et al., 2019), the bilinear nature of the reward in low-rank bandit problems makes the derivation of sampling lower bounds more challenging than for linear bandits. We overcome this challenge by considering a subset of perturbations of the reward matrix that behave linearly, yielding a relaxed sampling lower bound.
@@
Note that $M=PQ^{\top}$ yields $M_{i,j}=P_{i}^{\top}Q_{j}$ for every $(i,j)$. By using the same information-theoretic arguments as in the proof of the sampling lower bound for linear bandits and the previous characterization, we obtain the following result.
@@
$I=\inf_{\begin{subarray}{c}P^{\prime}\in\mathbb{R}^{m\times r},Q^{\prime}\in\mathbb{R}^{n\times r},\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2}\sum_{i,j}\omega_{i,j}(P_{i}^{\top}Q_{j}-P^{{}^{\prime}\top}_{i}Q^{\prime}_{j})^{2}.$
@@
$J:=\inf_{\begin{subarray}{c}M^{\prime},\operatorname{rank}(M^{\prime})=r,\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2}\sum_{i,j}\omega_{i,j}(M_{i,j}-M^{\prime}_{i,j})^{2}.$
@@
$J=\inf_{\begin{subarray}{c}M^{\prime},\operatorname{rank}(M^{\prime})\leq r,\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2}\sum_{i,j}\omega_{i,j}(M_{i,j}-M^{\prime}_{i,j})^{2}.$
@@
$J_{1}=\frac{2\varepsilon^{2}}{\sum_{i}\|Q^{i}_{\pi}\|^{2}_{\left(\Lambda^{i}_{Q}\right)^{-1}}}\text{ and }J_{2}=\frac{2\varepsilon^{2}}{\sum_{j}\|P^{j}_{\pi}\|^{2}_{\left(\Lambda^{j}_{P}\right)^{-1}}}.$
Proof of E.3.
@@
Consequently, $\omega_{\min}\Sigma\preccurlyeq\Lambda^{i}_{Q}\preccurlyeq\omega_{\max}\Sigma$ and $\omega_{\min}\Sigma\preccurlyeq\Lambda^{j}_{P}\preccurlyeq\omega_{\max}\Sigma$. In particular, since $\omega_{\min}>0,$ $\Lambda_{Q}^{i}$ and $\Lambda_{P}^{j}$ are invertible.
@@
$J_{1}=\inf_{\begin{subarray}{c}D\in\mathbb{R}^{m\times r},\\ \sum_{i}D_{i}^{\top}Q^{i}_{\pi}=2\varepsilon\end{subarray}}\frac{1}{2}\sum_{i}\|D_{i}\|^{2}_{\Lambda^{i}_{Q}},$
@@
$J_{2}=\inf_{\begin{subarray}{c}D\in\mathbb{R}^{n\times r},\\ \sum_{j}D_{j}^{\top}P^{j}_{\pi}=2\varepsilon\end{subarray}}\frac{1}{2}\sum_{j}\|D_{j}\|^{2}_{\Lambda^{j}_{P}}.$
@@
By combining Lemma E.2 and Lemma E.3, we directly get the main statement of Theorem 6.3. Finally, let us show that the lower bound is independent of the choice of rank factorization $M=PQ^{\top}$. Let $M=P^{\prime}Q^{\prime\top}$ be another rank factorization of $M$. Then, there exists an invertible matrix $R\in\mathbb{R}^{r\times r}$ such that $Q^{\prime}=QR$ and $P^{\prime}=PR^{-T}$ (Piziak & Odell, 1999). This implies that for every $(i,j)$, $Q^{\prime}_{j}=\left(Q_{j}^{\top}R\right)^{\top}=R^{\top}Q_{j},$ and $P^{\prime}_{i}=\left(P_{i}^{\top}R^{-T}\right)^{\top}=R^{-1}P_{i}$, which in turn ensures that $\Lambda^{i}_{Q^{\prime}}=R^{\top}\Lambda^{i}_{Q}R$ and $\Lambda^{j}_{P^{\prime}}=R^{\top}\Lambda^{j}_{P}R$ are invertible. Furthermore,
@@
$=\left(Q^{i}_{\pi}\right)^{\top}RR^{-1}\left(\Lambda^{i}_{Q}\right)^{-1}R^{-T}R^{T}Q^{i}_{\pi}$
@@
$T\geq\frac{\operatorname{kl}\left(\delta,1-\delta\right)}{2\sigma_{1}\omega_{\max}\varepsilon^{2}}\max\left(\sum_{i}\|Q^{i}_{\pi}\|_{2}^{2},\sum_{j}\|P^{j}_{\pi}\|_{2}^{2}\right).$
@@
By the proof of Lemma E.3, $\omega_{\min}\Sigma\preccurlyeq\Lambda^{i}_{Q}\preccurlyeq\omega_{\max}\Sigma$ and $\omega_{\min}\Sigma\preccurlyeq\Lambda^{j}_{P}\preccurlyeq\omega_{\max}\Sigma$. In particular, $\sigma_{r}\omega_{\min}\leq\lambda_{\min}\left(\Lambda^{i}_{Q}\right)\leq\lambda_{\max}\left(\Lambda^{i}_{Q}\right)\leq\sigma_{1}\omega_{\max}$, and the same results hold for $\Lambda^{j}_{P}$. Now, note that for any $x\in\mathbb{R}^{r}$ and any invertible matrix $B\in\mathbb{R}^{r\times r}$, $\lambda_{\min}\left(B\right)^{-1}\|x\|^{2}_{2}\geq\|x\|^{2}_{B^{-1}}\geq\lambda_{\max}\left(B\right)^{-1}\|x\|^{2}_{2}$. Consequently, for any $\left(i,j\right)$,

### v2

@@ C.3 Additional Lemmas
$\|\psi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\left(1+\frac{C}{2\sqrt{T}}\frac{L}{\|M\|_{\max}}\mu\kappa^{2}r^{1/2}\sqrt{\frac{mn}{\left(m+n\right)\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$
Combining this result with Proposition D.2 and noting that $\frac{\sigma L}{\|M\|_{\max}}\leq\frac{L^{2}}{\|M\|_{\max}}=A,$ we get that with probability larger than $1-\delta,$
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+C_{1}A\sqrt{\log(e/\delta)}\mu^{2}\kappa^{2}r\omega_{\min}^{-1}\left(m\wedge n\right)^{-1/2}\frac{\log^{3/2}\left(\left(m+n\right)T/\delta\right)}{T}$
$+C_{2}A\mu^{6}\kappa^{4}r^{3}\frac{\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
for some universal constants $C_{1},C_{2}>0$. Since $\sqrt{\log(e/\delta)}\mu^{2}\kappa^{2}r\omega_{\min}^{-1}\left(m\wedge n\right)^{-1/2}\log^{3/2}\left(\left(m+n\right)T/\delta\right)\lesssim\mu^{6}\kappa^{4}r^{3}\frac{\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\log^{3}\left(\left(m+n\right)T/\delta\right)$, we obtain that under the assumptions of Proposition D.2,
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\mu\sqrt{\frac{r\left(m+n\right)}{\omega_{\min}mn}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+K_{0}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$ (37)
with probability larger than $1-\delta,$ and transforming this result into a sample complexity guarantee as in the proof of Theorem 6.2 yields that $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt RS-PE}}|\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
The above sample complexity guarantee can be reformulated as
$T$ $\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)$
when $\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)\gtrsim\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$, i.e.,
$\varepsilon\lesssim\frac{\sigma^{2}\|M\|_{\max}}{L^{2}}\frac{\omega_{\min}mn\left(m\wedge n\right)}{\mu^{4}\kappa^{2}r^{2}\left(m+n\right)\sqrt{mn}}\frac{\log\left(e/\delta\right)}{\log^{3}\left(\left(m+n\right)T/\delta\right)}.$ (38)
When $M$ and $\omega$ are homogeneous and $\sigma\gtrsim\|M\|_{\max}$, this simplifies to
$T$ $\gtrsim\frac{\sigma^{2}\left(m+n\right)}{\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)$
@@
$\varepsilon\lesssim\frac{\|M\|_{\max}}{m+n}\frac{\log\left(e/\delta\right)}{\log^{3}\left(\left(m+n\right)T/\delta\right)}.$ (39)
@@ D.3 Additional Lemmas
A direct computation shows that the diagonal entries of $\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}$ can all be written as $\sum_{i,j}U^{2}_{i,k}V^{2}_{j,l}$, $\sum_{i,j}U^{2}_{\perp,i,k}V_{j,l}^{2}$ or $\sum_{i,j}U^{2}_{i,k}V_{\perp,j,l}^{2}$ for some $k$ and $l$. All of these terms are equal to $1$ by the orthogonality of $\begin{bmatrix}U&U_{\perp}\\ \end{bmatrix}$ and $\begin{bmatrix}V&V_{\perp}\\ \end{bmatrix}$. The entries outside of the diagonal can all be factorized by a scalar product between two columns of $\begin{bmatrix}U&U_{\perp}\\ \end{bmatrix}$ or two columns of $\begin{bmatrix}V&V_{\perp}\\ \end{bmatrix}$, which is equal to 0 by the same orthogonality argument. The same reasoning holds for $\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}$, as $\begin{bmatrix}\widehat{U}&\widehat{U}_{\perp}\\ \end{bmatrix}$ and $\begin{bmatrix}\widehat{V}&\widehat{V}_{\perp}\\ \end{bmatrix}$ are also orthogonal. The other part of the statement directly follows from the identities $\omega_{\max}\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}\succeq\Lambda\succeq\omega_{\min}\sum_{i,j}\psi_{i,j}\psi_{i,j}^{\top}$ and $\omega_{\max}\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}\succeq\Lambda_{\phi}\succeq\omega_{\min}\sum_{i,j}\phi_{i,j}\phi_{i,j}^{\top}$. ∎
@@
  • $\left|\|\phi_{\nu}\|^{2}_{2}-\|\psi_{\nu}\|^{2}_{2}\right|\leq 3\epsilon_{\textup{Sub-Rec}}$
@@
$\phi_{i,j}^{\top}\phi_{i^{\prime},j^{\prime}}=\left(\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}\right)\left(\widehat{V}_{j}^{\top}\widehat{V}_{j^{\prime}}\right)+\left(\widehat{U}_{\perp,i}^{\top}\widehat{U}_{\perp,i^{\prime}}\right)\left(\widehat{V}_{j}^{\top}\widehat{V}_{j^{\prime}}\right)+\left(\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}\right)\left(\widehat{V}_{\perp,j}^{\top}\widehat{V}_{\perp,j^{\prime}}\right)$
@@
By orthogonality arguments, $\widehat{U}_{i}^{\top}\widehat{U}_{i^{\prime}}+\widehat{U}_{\perp,i}^{\top}\widehat{U}_{\perp,i^{\prime}}=U_{i}^{\top}U_{i^{\prime}}+U_{\perp,i}^{\top}U_{\perp,i^{\prime}}=0$ when $i\neq i^{\prime}$, and $\widehat{V}_{j}^{\top}\widehat{V}_{j^{\prime}}+\widehat{V}_{\perp,j}^{\top}\widehat{V}_{\perp,j^{\prime}}=V_{j}^{\top}V_{j^{\prime}}+V_{\perp,j}^{\top}V_{\perp,j^{\prime}}=0$ when $j\neq j^{\prime}$. Thus, two out of the three terms in the equalities above cancel out when $(i,j)\neq(i^{\prime},j^{\prime})$. When $(i,j)=(i^{\prime},j^{\prime})$, we can combine two terms by using the fact that $\|U_{i}\|^{2}_{2}+\|U_{\perp,i}\|^{2}_{2}=1$ to write $\psi_{i,j}^{\top}\psi_{i,j}=\|V_{j}\|^{2}_{2}+\|U_{i}\|^{2}_{2}\|V_{\perp,j}\|^{2}_{2}.$ In any case, each term in the difference can be controlled by the subspace recovery error. For example,
@@
$\leq 2\epsilon_{\textup{Sub-Rec}}.$
@@
$\phi_{i,j}^{\top}\phi_{i^{\prime},j^{\prime}}-\psi_{i,j}^{\top}\psi_{i^{\prime},j^{\prime}}\leq\left(2+\mathds{1}_{(i,j)=(i^{\prime},j^{\prime})}\right)\epsilon_{\textup{Sub-Rec}}.$
@@
$\leq\left(\sum_{(i,j),(i^{\prime},j^{\prime})}\nu_{i,j}\nu_{i^{\prime},j^{\prime}}\left(2+\mathds{1}_{(i,j)=(i^{\prime},j^{\prime})}\right)\right)\epsilon_{\textup{Sub-Rec}}$
$=\left(2+\sum_{i,j}\nu_{i,j}^{2}\right)\epsilon_{\textup{Sub-Rec}},$
@@
There are two interesting things to mention about the second inequality of Lemma D.12. First, it is tight in the sense that it is attained for $r=m\wedge n$. Indeed, assume then that $m\wedge n=n$ without loss of generality. On the one hand, $\|\psi_{i,j}\|^{2}_{2}=\|V_{j}\|^{2}_{2}+\|U_{i}\|^{2}_{2}\left(1-\|V_{j}\|^{2}_{2}\right)=1$ since $V_{j}$ is a row of the orthogonal matrix $V$. On the other hand, by orthogonality of $U$, $\mu=\|U\|_{2\to\infty}=1$ so that $\frac{\mu^{2}r\left(m+n-\mu^{2}r\right)}{mn}=1$. Second, it directly entails an upper bound on the norm of $\psi_{\pi}.$ Indeed,
$\|\psi_{\pi}\|_{2}=\|\sum_{i,j}\omega^{\pi}_{i,j}\psi_{i,j}\|_{2}\leq\sum_{i,j}\omega^{\pi}_{i,j}\|\psi_{i,j}\|_{2}\leq\max_{i,j}\|\psi_{i,j}\|_{2},$
where we have used $\sum_{i,j}\omega^{\pi}_{i,j}=1$ in the last inequality. This ensures that
@@
In this appendix, we assume that $\xi_{t}\sim\mathcal{N}(0,\sigma^{2})$ and we choose the rank factorization $M=U\Sigma V^{\top}=PQ^{\top}$ with $P=U\sqrt{\Sigma}$ and $Q=V\sqrt{\Sigma}$. Furthermore, we denote by $N_{i}$ the $i$-th row vector of a matrix $N$. As noted in (Jun et al., 2019), the bilinear nature of the reward in low-rank bandit problems makes the derivation of sampling lower bounds more challenging than for linear bandits. We overcome this challenge by considering a subset of perturbations of the reward matrix that behave linearly, yielding a relaxed sampling lower bound.
@@
Note that $M=PQ^{\top}$ yields $M_{i,j}=P_{i}^{\top}Q_{j}$ for every $(i,j)$. By using the same arguments as in the proof of the sampling lower bound for linear bandits (see Proposition C.2) and the characterization of Lemma E.1, we obtain the following result.
@@
$I=\inf_{\begin{subarray}{c}P^{\prime}\in\mathbb{R}^{m\times r},Q^{\prime}\in\mathbb{R}^{n\times r},\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2\sigma^{2}}\sum_{i,j}\omega_{i,j}(P_{i}^{\top}Q_{j}-P^{{}^{\prime}\top}_{i}Q^{\prime}_{j})^{2}.$
@@
$J:=\inf_{\begin{subarray}{c}M^{\prime},\operatorname{rank}(M^{\prime})=r,\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2\sigma^{2}}\sum_{i,j}\omega_{i,j}(M_{i,j}-M^{\prime}_{i,j})^{2}.$
@@
$J=\inf_{\begin{subarray}{c}M^{\prime},\operatorname{rank}(M^{\prime})\leq r,\\ |v^{\pi}-v^{{}^{\prime}\pi}|\geq 2\varepsilon\end{subarray}}\frac{1}{2\sigma^{2}}\sum_{i,j}\omega_{i,j}(M_{i,j}-M^{\prime}_{i,j})^{2}.$
@@
$J_{1}=\frac{2\varepsilon^{2}}{\sigma^{2}\sum_{i}\|Q^{i}_{\pi}\|^{2}_{\left(\Lambda^{i}_{Q}\right)^{-1}}}\text{ and }J_{2}=\frac{2\varepsilon^{2}}{\sigma^{2}\sum_{j}\|P^{j}_{\pi}\|^{2}_{\left(\Lambda^{j}_{P}\right)^{-1}}}.$
Proof of Lemma E.3.
@@
Consequently, $\omega_{\min}\Sigma\preceq\Lambda^{i}_{Q}\preceq\omega_{\max}\Sigma$ and $\omega_{\min}\Sigma\preceq\Lambda^{j}_{P}\preceq\omega_{\max}\Sigma$. In particular, since $\omega_{\min}>0,$ $\Lambda_{Q}^{i}$ and $\Lambda_{P}^{j}$ are invertible.
@@
$J_{1}=\inf_{\begin{subarray}{c}D\in\mathbb{R}^{m\times r},\\ \sum_{i}D_{i}^{\top}Q^{i}_{\pi}=2\varepsilon\end{subarray}}\frac{1}{2\sigma^{2}}\sum_{i}\|D_{i}\|^{2}_{\Lambda^{i}_{Q}},$
@@
$J_{2}=\inf_{\begin{subarray}{c}D\in\mathbb{R}^{n\times r},\\ \sum_{j}D_{j}^{\top}P^{j}_{\pi}=2\varepsilon\end{subarray}}\frac{1}{2\sigma^{2}}\sum_{j}\|D_{j}\|^{2}_{\Lambda^{j}_{P}}.$
@@
By combining Lemmas E.2 and E.3, we directly get the main statement of Theorem 6.3. Finally, let us show that the lower bound is independent of the choice of rank factorization $M=PQ^{\top}$. Let $M=P^{\prime}Q^{\prime\top}$ be another rank factorization of $M$. Then, there exists an invertible matrix $R\in\mathbb{R}^{r\times r}$ such that $Q^{\prime}=QR$ and $P^{\prime}=PR^{-\top}$ (Piziak & Odell, 1999). This implies that for every $(i,j)$, $Q^{\prime}_{j}=\left(Q_{j}^{\top}R\right)^{\top}=R^{\top}Q_{j},$ and $P^{\prime}_{i}=\left(P_{i}^{\top}R^{-\top}\right)^{\top}=R^{-1}P_{i}$, which in turn ensures that $\Lambda^{i}_{Q^{\prime}}=R^{\top}\Lambda^{i}_{Q}R$ and $\Lambda^{j}_{P^{\prime}}=R^{\top}\Lambda^{j}_{P}R$ are invertible. Furthermore,
@@
$=\left(Q^{i}_{\pi}\right)^{\top}RR^{-1}\left(\Lambda^{i}_{Q}\right)^{-1}R^{-\top}R^{\top}Q^{i}_{\pi}$
@@
$T\geq\frac{\sigma^{2}\operatorname{kl}\left(\delta,1-\delta\right)}{2\sigma_{1}\omega_{\max}\varepsilon^{2}}\max\left(\sum_{i}\|Q^{i}_{\pi}\|_{2}^{2},\sum_{j}\|P^{j}_{\pi}\|_{2}^{2}\right).$
@@
By the proof of Lemma E.3, $\omega_{\min}\Sigma\preceq\Lambda^{i}_{Q}\preceq\omega_{\max}\Sigma$ and $\omega_{\min}\Sigma\preceq\Lambda^{j}_{P}\preceq\omega_{\max}\Sigma$. In particular, $\sigma_{r}\omega_{\min}\leq\lambda_{\min}\left(\Lambda^{i}_{Q}\right)\leq\lambda_{\max}\left(\Lambda^{i}_{Q}\right)\leq\sigma_{1}\omega_{\max}$, and the same results hold for $\Lambda^{j}_{P}$. Now, note that for any $x\in\mathbb{R}^{r}$ and any invertible matrix $B\in\mathbb{R}^{r\times r}$, $\lambda_{\min}\left(B\right)^{-1}\|x\|^{2}_{2}\geq\|x\|^{2}_{B^{-1}}\geq\lambda_{\max}\left(B\right)^{-1}\|x\|^{2}_{2}$. Consequently, for any $\left(i,j\right)$,